# Look Before you Leap: Estimating LLM Benchmark Scores from Descriptions

### v2

@@ Abstract
Progress in large language models is constrained by an evaluation bottleneck: build a benchmark, run models, then iterate. We ask a question: can we forecast outcomes before running any experiments to inform earlier study design? For example, a team building an AI assistant for a certain task can estimate whether expected performance is around 50 or closer to 80, evidence that supports whether to proceed to a pilot study, how to scope it, and how to allocate resources. We study text-only performance forecasting, where a model predicts a score from a redacted task description and intended configuration, with no access to dataset instances. To support systematic study, we curate Precog, a corpus of redacted description–performance pairs spanning diverse tasks, domains, and metrics. We scrape task and configuration descriptions from arXiv, yielding 2,290 instances covering 1,519 papers, and construct a leakage-free test split using papers published after the evaluated models’ knowledge cutoff. Experiments show the task is challenging but feasible: reasoning models achieve moderate prediction performance with well-calibrated uncertainty, reaching mean absolute error as low as $9.9$ at high-confidence thresholds. We further test a zero-leakage setting—forecasting on newly released datasets or experiments before their papers are indexed—where GPT-5 with built-in web search still attains non-trivial prediction accuracy. Overall, our corpus and analyses offer an initial step toward open-ended anticipatory evaluation, supporting difficulty estimation and smarter experiment prioritization. [^1] Our data and code and are available in https://github.com/JJumSSu/PRECOG
@@ 1 Introduction
Modern LLM development is evaluation-centric: define a benchmark, evaluate models and configurations, iterate (brown2020language; achiam2023gpt; touvron2023llama; team2023gemini). As benchmarks proliferate, exhaustive curation and full experiments demand substantial compute, researcher time, and funding (mazumder2023dataperf). We therefore seek an ex ante performance forecaster to help designers decide what to build and where to measure. For example, consider a team proposing a binary “solver-choice” task on an ordinary differential equations benchmark (Figure 1; gaonkar2025sciml). Before annotating a single example, they can ask a concrete what-if: will GPT-4.1 (achiam2023gpt) without CoT (wei2022chain) land nearer 30 or 60 Accuracy points? A forecast like this has the potential to quickly inform scope and spend—how large a pilot to run, which baselines to include, and whether the task merits full-scale annotation.
Figure 1: Precog pairs a redacted task/dataset description with its target metric. Given the inputs (a redacted description of dataset and an experimental configuration (gaonkar2025sciml)), the goal is to predict the target metric. The regressor conditions on performance-relevant details and may retrieve corroborating evidence from the literature.
Two adjacent lines of work on ML model performance prediction address related goals: (i) missing-entry prediction, which infers unobserved cells from observed ones (xia2020predicting; ye2021finegrained; ye2023predictable; zhaocan); and (ii) efficient evaluation, which extrapolates full-benchmark performance from a small set of pilot runs (maiapolo2024tinybenchmarks; pacchiardi2024hundred; anugraha2025proxylm). Both assume seed measurements or closely related tasks and typically rely on engineered numeric features. In contrast, we target a zero-example, open-ended setting in which retrieval is driven by natural-language task descriptions, supporting researchers who are creating new datasets or designing new experiments. Prior work has also explored the capability of LLMs to forecast the outcomes of future geopolitical events in online prediction markets (kargerforecastbench; halawi2024approaching), however to the best of our knowledge, this is the first work to explore the capability of LLMs to predict their own performance on a new task based only on a text description.
@@
We formulate text-only performance forecasting as regression: given a natural-language specification of the dataset and evaluation setup, including task, data provenance, prompting/decoding choices, and metric/protocol, we predict a target metric normalized to a 0100100 scale. As in Figure 1, the input is a description of the proposed task and configuration; the output is a predicted metric value (e.g., $32$). Unlike feature-engineered baselines, we operate directly on free-form, literature-style text and optionally use retrieval to consult related prior work, rather than interpolation from seed results.
To enable systematic study, we curate Precog, an arXiv-mined corpus of experimental records (park2025can) that, for each entry, links the benchmark’s source paper (introducing the dataset) to the paper reporting results on it. From these papers we extract task and configuration descriptions, yielding 2,290 instances spanning 1,519 papers, and we construct a leakage-free test split by restricting to papers published after the evaluated models’ knowledge cutoff (2025 January-September). From these sources, we synthesize redacted descriptions that preserve task and setup details while masking dataset names and numeric results. To further probe residual leakage risks (tian2024scicode), we also evaluate a streaming regime that ingests arXiv releases and makes predictions before they are indexed by search engines.
We then evaluate zero-shot LLM regressors, optionally augmented with literature search. Given a redacted description, the model may iteratively issue evidence-seeking queries (ReAct-style (yao2023react)) and outputs a point estimate alongside a self-reported confidence (yoon2025reasoning). We assess calibration by sweeping a confidence threshold and reporting the coverage–MAE trade-off, and we analyze search tool-use behavior.
On Precog, a proprietary reasoning model attains an MAE of $14.6$ across seven metric groups, and $9.9$ on a high-confidence, accuracy-filtered subset, with reasonably calibrated uncertainty. We perform both reasoning and subset ablations to identify which factors drive performance prediction and to characterize its behavior. In the streaming setting, GPT-5 (openai2025gpt5) with built-in web search delivers competitive, though imperfect, performance on newly released datasets, while maintaining good calibration.
Altogether, our study shows that LLM-based regressors that predict model performance from text descriptions can inform benchmark design, prioritize configurations, determine dataset sizes, and plan pilot experiments, while reducing compute and turnaround time for exploratory analysis. We release Precog, a redacted-description benchmark grounded in the scientific literature, together with prompts, protocols, and code. We hope these resources and our evaluation methodology will support future advances in predictive evaluation and model selection.
@@ 2 Related Work
Work Text. Feat. Runs Idea Task Reg.
xia2020predicting
ye2021finegrained
ye2023predictable
maiapolo2024tinybenchmarks
pacchiardi2024hundred
anugraha2025proxylm
wen2025predicting
@@
Table 1: Comparison of performance-prediction schemas. Text.: task descriptions; Feat.: numeric features; Runs: running the model on a subset of the benchmark; Idea: predicts unseen method’s performance; Task: predicts performance on an unseen task; Reg.: (Regression) predicts an metric value.
Prior works primarily fall into two categories. (1) Missing-entry prediction: infer unobserved cells from observed ones using matrix-completion or response-surface models (achille2019task2vec; xia2020predicting; chen2021model; ye2021finegrained; ye2023predictable; rabinovich2023predicting; shi2025importance; zhaocan); these assume seed results, stable task identities, and overlap for interpolation. (2) Efficient evaluation: run a small subset of trials and extrapolate full-benchmark performance (maiapolo2024tinybenchmarks; pacchiardi2024hundred; anugraha2025proxylm); these still require executing target (or closely related) runs. Finally, we note contemporaneous work by wen2025predicting, which aims to predict which of two novel research ideas is likely to perform better on a set of existing benchmarks. In contrast, our work demonstrates that LLMs can predict the absolute performance of a fixed model on a novel benchmark. See Table 1 for a detailed comparison.
LLMs for Regression
LLMs have been used as regressors in structured settings (AutoML/black-box optimization) (song2024omnipred; vacareanu2024words; nguyen2024predicting; requeima2024llm; tang2024understanding), often relying on key–value hyperparameter strings and access to seed trials or finetuning on the target task. In contrast, we study a zero-run, description-only regression problem where the sole input is a free-form specification of a novel dataset and evaluation protocol. Our task involves forecasting, specifically, predicting the outcomes of ideas before conducting real experiments. (zou2022forecasting; schoenegger2023large; kargerforecastbench; abolghasemi2025humans; moussa2025scholareval). Inspired by these works, our work demonstrates the potential of reasoning LLMs in forecasting the outcomes of evaluation results on unseen datasets.
Extracting Structured Records from Litertaure
Prior work extracts structured result tuples (e.g., task, dataset, metric) from scientific articles (singh2019automated; hou2019identification; kardas2020axcell; yang2022telin; bai2023schema; csahinucc2024efficient; park2025can). Precog builds on these tuples: for each experimental record, we synthesize a richer natural-language description that augments the tuple with additional context.
@@ 3 Task
We study zero-example, text-only performance forecasting as a regression problem. Given a self-contained, redacted description $x_{i}$ of a dataset and evaluation configuration, the model performs reasoning and outputs a numeric estimate $\hat{y}_{i}\in[0,100]$. We treat $x_{i}$ as a structured yet free-form text description that captures contextual nuances of the task, dataset, and experimental configuration without requiring access to instances. To compare heterogeneous metrics (e.g., Accuracy, F1, BLEU), we convert each paper-reported score $y_{i}^{\text{paper}}$ to a common 0–100 scale via a simple normalization $(N:\mathbb{R}\!\to\![0,100])$; $y_{i}=N(y_{i}^{\text{paper}})$. Concretely, values already on a 0–100 scale remain unchanged, while fractional scores in $([0,1])$ are multiplied by 100 (e.g., Accuracy $(0.35\mapsto 35.0)$; F1 $(83.2\mapsto 83.2)$.
Given a redacted description $x_{i}$, an LLM regressor $f_{\theta}$ predicts the normalized target as $\hat{y}_{i}=f_{\theta}(x_{i},Z_{i})$, where $Z_{i}$ is an optional set of retrieved evidence. We obtain $Z_{i}$ by querying a time-bounded literature corpus $\mathcal{C}$ with a top-$k$ retriever $R_{k}$ on a query formed from $x_{i}$, i.e., $Z_{i}=R_{k}(q(x_{i}),\,\mathcal{C})$. Setting $k=0$ yields the description-only variant $f_{\theta}(x_{i},\varnothing)$. At inference, $f_{\theta}$ may iterate in a ReAct-style loop—reason (yao2023react), issue $q(\cdot)$, retrieve $Z_{i}$, refine—and then output $\hat{y}_{i}$. For example, for the open-domain QA description above, $f_{\theta}$ can retrieve prior exact-match (EM) ranges for similar setups and use them to refine $\hat{y}_{i}$.
@@
Following prior work (srinivasan2021predicting; ahuja2022multi; jawahar2024llmpp), we evaluate $f_{\theta}$ with two standard criteria: mean absolute error (MAE) and the Pearson correlation $\mathrm{r}$ between $y_{i}$ and $\hat{y}_{i}$. MAE (lower is better) measures absolute accuracy across heterogeneous metrics, while $\mathrm{r}$ (higher is better) assesses whether predictions preserve the relative ordering of tasks/configurations. Used together, they capture both value accuracy and trend fidelity despite differences in metric scales and variances—an evaluation protocol that is standard for performance prediction.
@@ 4 Data Curation
To evaluate at scale, we need a large number of diverse and realistic, open-ended pairs $(x_{i},y_{i})$. Rather than synthesizing experimental configurations and running automated experiments, which is both costly and unlikely to capture the kinds of problems practitioners actually investigate, we automatically curate experimental configurations and their reported results directly from the scientific literature. Our pipeline harvests redacted task descriptions and reported scores from published papers. We follow the pipeline in Figure 2: (i) collect arXiv experimental records (section 4.1); (ii) define the record provenance and paper types (section 4.2); (iii) recover each record’s sources (section 4.3); (iv) extract a schema-aligned, redacted description (section 4.4); and (v) perform validation and quality control (section 4.5).
4.1 Source Corpus
@@
We build on LLMEvalDB (park2025can), focusing on its 2023–2024 entries and 2025 entries from January to September, which are dominated by non-finetunable, proprietary models. This choice reduces variance from model/method tuning and better isolates dataset/task effects. Moreover, because current LLMs generally have a 2024 knowledge cutoff, restricting the 2025 split to post-cutoff papers mitigates leakage: the reported metrics could not have appeared in the training data. Each experimental record is a table cell with the dataset name, configuration context, and reported metric. Values are normalized to define $y_{i}$. We consider seven metrics with natural [0,100] scales: Accuracy, F1, Recall, Precision, ROUGE, BLEU, and EM.
4.2 Record Provenance and Paper Types
@@
Each record is grounded in two literature sources: the result paper, which reports the score, and the dataset paper, which introduced the dataset. These may coincide or differ. For example, Paper A introduces a QA dataset and reports accuracy for several models (result paper = dataset paper). Later, Paper B evaluates a new prompting/evaluation protocol on the same dataset under a revised setup (result paper $\neq$ dataset paper) (park2021faviq; rangapur2024battle).
4.3 Dataset Source Paper Retrieval
@@
To compile faithful descriptions for each record, we first collect its sources. LLMEvalDB (park2025can) provides the result papers. When the result paper differs from the dataset paper, we automatically recover the latter by inferring the citation with an LLM, resolving it to an arXiv ID, downloading the PDF, and verifying the dataset name/title. We then run a rule-based check to confirm that the PDF explicitly mentions the dataset (string match).
4.4 Extracting Redacted Input Description
@@
Given the sources for the result and dataset papers, an extractor LLM produces a schema-aligned, redacted description $x_{i}$ that summarizes the task (problem definition and input/label format), data collection (provenance, domain, collection protocol), evaluation (metric and protocol), difficulty cues (number of classes, long context, reasoning depth), the relevant subset/split, the prompting strategy (zero/few-shot, CoT), and any other record-specific settings (context-window, tool/retrieval use), when reported; see the Appendix E for illustrative examples.
@@
When the retrieved sources do not actually cover the experimental record (e.g., due to retrieval errors from 4.3), we discard the case by having the extractor return failed, thereby correcting residual retrieval mistakes. For the remaining records, we remove dataset names, numerical results, and summary statistics to produce faithful, self-contained, anonymized.
Figure 2: Overview of the Precog curation pipeline. For each record, we combine the result paper (mitchell2023comparing), which reports the score and specifies the evaluation protocol and configuration, with the dataset paper (moskvichev2023conceptarc), which provides the task description, to produce a schema-aligned input. The score $0.30$ reported is normalized to serve as the target.
4.5 Data Statistics and Quality Control
@@
We construct a bounded yet diverse corpus via stratified sampling over source papers, enforcing caps per paper and per dataset plus per-source quotas to prevent any single source from dominating. This yields 2,290 instances of records spanning distinct 1,019 dataset papers and 1,022 unique result papers, covering many datasets and configurations rather than a single benchmark (ye2023predictable). Because Precog is LLM-assisted, we further audited 30 random descriptions under a fixed protocol—checking anonymization (binary: no dataset names or performance numbers), schema coverage (5-point), and source support vs. the PDFs (5-point). Results (Table 2) show well-anonymized, accurate, and comprehensively grounded descriptions, suitable as schema-aligned, leakage-resistant inputs for text-only performance forecasting.

### v3

# Anticipatory Evaluation of Language Models

@@ Abstract
Progress in large language models is increasingly constrained by an evaluation bottleneck: benchmarks must be built and models run before iteration can begin. We investigate whether evaluation outcomes can be forecast before any experiments are conducted. Specifically, we study text-only performance prediction, where models estimate performance from task descriptions and experimental configurations alone, without access to dataset instances. To support systematic study, we curate Precog, a corpus of description–performance pairs spanning diverse tasks, domains, and metrics. We scrape task and configuration descriptions from arXiv, yielding 2,290 instances covering 1,519 papers, and construct a test split using papers published after the evaluated models’ knowledge cutoff. Experiments show the task is challenging but feasible: reasoning models achieve a non-trivial forecasting skill reaching mean absolute error as low as $9.9$ at high-confidence thresholds. Overall, our corpus and analyses offer an initial step toward open-ended anticipatory evaluation, supporting difficulty estimation and smarter resource allocation.
Machine Learning, ICML, performance prediction, regression, AI4Science, LLM
@@ 1 Introduction
Performance prediction has long been a central problem in machine learning research. Prior work has studied how to forecast model performance across datasets, tasks, and configurations, including predicting missing experimental results from partially observed tables (Xia et al., 2020; Ye et al., 2021, 2023; Zhao et al., 2025), estimating full-benchmark performance from small pilot runs (Maia Polo et al., 2024; Pacchiardi et al., 2024; Anugraha et al., 2025), and modeling trends across model scale, data, and compute (Kaplan et al., 2020; Hoffmann et al., 2022). Related efforts in AutoML and black-box optimization also rely on performance prediction to guide model and hyperparameter search, often treating evaluation outcomes as a regression problem over experimental configurations (Song et al., 2024; Requeima et al., 2024). These approaches have become increasingly important as modern evaluation grows more expensive and complex.
However, existing performance prediction methods largely operate after experiments have begun. They typically assume access to seed measurements, closely related datasets, structured experimental tables, or pilot evaluations. In contrast, many of the most consequential design decisions occur before any data instances exist: deciding whether a benchmark is likely to be too easy or too hard, which experimental configurations are worth labeling data to support, or whether a proposed evaluation will meaningfully differentiate model classes. This paper studies a challenging but practically important problem: forecasting model performance before any experiments are run. Specifically, we aim to estimate expected evaluation outcomes using only a natural-language description of a proposed task and experimental setup, prior to annotation, benchmarking, or pilot studies.
Figure 1: Precog pairs a benchmark design brief (task description) with a target metric. Rather than running a pilot that requires data labeling and model inference, description-based performance prediction can make pilot studies much more efficient. Given (i) a textual specification of the benchmark (e.g., the number of multiple-choice options and how distractors are constructed) and (ii) an experimental configuration (e.g., greedy decoding with no chain-of-thought), the goal is to predict performance on the target metric. The regressor focuses on performance-relevant details and can leverage results from existing, related benchmarks.
While ambitious, we hypothesize that such ex ante performance estimates could play a valuable role in early-stage experiment planning. Textual benchmark descriptions often encode regularities about task structure, data provenance, evaluation protocols, and known sources of difficulty. If models can exploit these signals, even imperfect forecasts may help researchers make better decisions earlier in the planning phase. For example, a team designing a multiple-choice QA benchmark might wish to quantify how the number of answer options affects difficulty (see Figure 1). Before committing resources to large-scale annotation, they could pose a concrete what-if question derived directly from a draft problem specification: how would increasing the number of choices (e.g., $5\rightarrow 10\rightarrow 20$) affect performance for a target class of models? Linking design decisions to expected difficulty could inform pilot experiment design, baseline selection, and provide an early estimate as to whether the benchmark is likely to be discriminative at all.
@@
We formalize this setting as text-only performance forecasting. Unlike prior work on model performance prediction, we assume zero observed experimental results: given only a natural-language specification of the dataset and evaluation setup including the task, data provenance, prompting and decoding choices, and the evaluation protocol, we predict a target metric normalized to a 0100100 scale.
Our experiments demonstrate that text-only performance forecasting is challenging but feasible with current frontier models. Across 2,290 benchmark-configuration pairs, GPT-5 achieves a mean absolute error (MAE) of $14.6$ on a leakage-controlled 2025 test split, substantially outperforming a strong text-embedding baseline. When restricting to high-confidence predictions, error drops further to $9.9$ MAE, indicating that self-reported confidence provides a useful signal for identifying reliable forecasts. Performance remains stable across pre-knowledge cutoff (2023–2024) and post-cutoff (2025) splits, suggesting that prediction accuracy is not driven by memorization or contamination. In a fully streaming setting, forecasting results from newly released arXiv papers collected before search indexing, GPT-5 achieves $14.0$ MAE and $0.78$ Pearson correlation, closely matching offline performance and supporting the feasibility of real-time anticipatory evaluation. Finally, in a human comparison study, expert forecasters achieve $19.6$ MAE, while GPT-5 attains $13.6$ MAE on the same subset, underscoring the potential value of LLM-based predictors as early-stage decision aids.
Overall, reasoning LLMs can serve as a lightweight planning tool for zero-result settings by predicting difficulty upfront and prioritizing configurations. We release Precog, a literature-grounded benchmark of task and experiment descriptions, with prompts and code to support predictive evaluation and more efficient experimental design. [^1] Our data and code are available in https://github.com/JJumSSu/PRECOG
@@ 2 Related Work
Prior works primarily fall into two categories. (1) Missing-entry prediction: infer unobserved cells from observed ones using matrix-completion or response-surface models (Achille et al., 2019; Xia et al., 2020; Chen and Ritter, 2021; Ye et al., 2021, 2023; Rabinovich et al., 2023; Shi et al., 2025; Zhao et al., 2025); these assume seed results, stable task identities, and overlap for interpolation. (2) Efficient evaluation: run a small subset of trials and extrapolate full-benchmark performance (Maia Polo et al., 2024; Pacchiardi et al., 2024; Anugraha et al., 2025); these still require executing target (or closely related) runs. Finally, we note contemporaneous work by Wen et al. (2025), which aims to predict which of two novel research ideas is likely to perform better on a set of existing benchmarks. In contrast, our work demonstrates that LLMs can predict the absolute performance of a fixed model on a novel benchmark. See Table 1 for a detailed comparison.
LLMs for Regression & Forecasting
LLMs have been used as regressors in structured settings (AutoML/black-box optimization) (Song et al., 2024; Vacareanu et al., 2024; Nguyen et al., 2024; Requeima et al., 2024; Tang et al., 2024), often relying on key–value hyperparameter strings and access to seed trials or finetuning on the target task. In contrast, we study a zero-run, description-only regression problem where the sole input is a free-form specification of a novel dataset and evaluation protocol. Our task is a forecasting problem: we aim to predict the outcomes of ideas before running real experiments.(Zou et al., 2022; Schoenegger and Park, 2023; Karger et al., 2024; Halawi et al., 2024; Abolghasemi et al., 2025; Moussa et al., 2025). A growing body of work has investigated LLM-based approaches to time-series forecasting (Gruver et al., 2023; Jin et al., 2023; Cao et al., 2023; Ansari et al., 2024; Das et al., 2024; Tan et al., 2024; Ning et al., 2025; Jiang et al., 2025). Inspired by these works, our work demonstrates the potential of reasoning LLMs in forecasting the outcomes of evaluation results on unseen datasets.
Extracting Structured Records from Litertaure
Prior work extracts structured result tuples (e.g., task, dataset, metric) from scientific articles (Singh et al., 2019; Hou et al., 2019; Kardas et al., 2020; Yang et al., 2022; Bai et al., 2023; Şahinuç et al., 2024; Park et al., 2025). Precog builds on these tuples: for each experimental record, we synthesize a richer natural-language description that augments the tuple with additional context.
Table 1: Comparison of performance-prediction schemas. Text: uses text descriptions as input; Task: predicts performance on unseen tasks; Reg.: predicts a metric value (regression); No Feat.: requires no manual feature engineering; No Runs: does not require running the model on the benchmark.
@@
Work Text Task Reg. No Feat. No Runs
Xia et al. (2020)
Ye et al. (2021)
Ye et al. (2023)
Maia Polo et al. (2024)
Pacchiardi et al. (2024)
Anugraha et al. (2025)
Wen et al. (2025)
@@
Figure 2: Overview of the Precog curation pipeline. For each experimental record, we pair an evaluation/result paper (Mitchell et al., 2023), which reports the score and details the evaluation protocol and configuration, with the corresponding dataset paper (Moskvichev et al., 2023), which describes the task. We then map these sources into a single schema-aligned input. The reported score of $0.30$ is converted to a 0100100 scale (i.e., $0.30\mapsto 30$) and used as the prediction target.
@@ 3 Task
We study zero-example, text-only performance forecasting as a regression problem. Given a self-contained, identity-masked description $x_{i}$ of a dataset and evaluation configuration, the model performs reasoning and outputs a numeric estimate $\hat{y}_{i}\in[0,100]$. We treat $x_{i}$ as a structured yet free-form text description that captures contextual nuances of the task, dataset, and experimental configuration without requiring access to instances. To compare heterogeneous metrics (e.g., Accuracy, F1, BLEU), we convert each paper-reported score $y_{i}^{\text{paper}}$ to a common 0–100 scale via a simple normalization $(N:\mathbb{R}\!\to\![0,100])$; $y_{i}=N(y_{i}^{\text{paper}})$. Concretely, values already on a 0–100 scale remain unchanged, while fractional scores in $([0,1])$ are multiplied by 100 (e.g., Accuracy $(0.35\mapsto 35.0)$; F1 $(83.2\mapsto 83.2)$.
Given a description $x_{i}$, an LLM regressor $f_{\theta}$ predicts the normalized target as $\hat{y}_{i}=f_{\theta}(x_{i},Z_{i})$, where $Z_{i}$ is an optional set of retrieved evidence. We obtain $Z_{i}$ by querying a time-bounded literature corpus $\mathcal{C}$ with a top-$k$ retriever $R_{k}$ on a query formed from $x_{i}$, i.e., $Z_{i}=R_{k}(q(x_{i}),\,\mathcal{C})$. Setting $k=0$ yields the description-only variant $f_{\theta}(x_{i},\varnothing)$. At inference, $f_{\theta}$ may iterate in a ReAct-style loop—reason (Yao et al., 2023), issue $q(\cdot)$, retrieve $Z_{i}$, refine—and then output $\hat{y}_{i}$. For example, for the open-domain QA description above, $f_{\theta}$ can retrieve prior exact-match (EM) ranges for similar setups and use them to refine $\hat{y}_{i}$.
@@
Following prior work (Srinivasan et al., 2021; Ahuja et al., 2022; Jawahar et al., 2024), we evaluate $f_{\theta}$ with two standard criteria: mean absolute error (MAE) and the Pearson correlation $\mathrm{r}$ between $y_{i}$ and $\hat{y}_{i}$. MAE (lower is better) measures absolute accuracy across heterogeneous metrics, while $\mathrm{r}$ (higher is better) assesses whether predictions preserve the relative ordering. Used together, they capture both value accuracy and trend fidelity despite differences in metric scales and variances—an evaluation protocol that is standard for performance prediction.
@@ 4 Data Curation
To evaluate at scale, we need a large number of diverse and realistic, open-ended pairs $(x_{i},y_{i})$. Rather than synthesizing experimental configurations and running automated experiments, which is both costly and unlikely to capture the kinds of problems practitioners actually investigate, we automatically curate experimental configurations and their reported results directly from the scientific literature. Our pipeline harvests task descriptions and reported scores from published papers. We follow the pipeline in Figure 2: (i) collect arXiv experimental records; (ii) define the record provenance and paper types; (iii) recover each record’s sources; (iv) extract a schema-aligned, identity-masked description; and perform validation and quality control.
Source Corpus
@@
We build on LLMEvalDB (Park et al., 2025), focusing on its 2023–2024 entries and 2025 entries from January to September, which are dominated by non-finetunable, proprietary models. This choice reduces variance from model/method tuning and better isolates dataset/task effects. Moreover, because current LLMs generally have a 2024 knowledge cutoff, restricting the 2025 split to post-cutoff papers mitigates leakage: the reported metrics could not have appeared in the training data. Each experimental record is a table cell with the dataset name, configuration context, and reported metric. Values are normalized to define $y_{i}$. We consider seven metrics with natural [0,100] scales: Accuracy, F1, Recall, Precision, ROUGE, BLEU, and EM.
Record Provenance and Paper Types
@@
Each record is grounded in two literature sources: the result paper, which reports the score, and the dataset paper, which introduced the dataset. These may coincide or differ. For example, Paper A introduces a QA dataset and reports accuracy for several models (result paper=dataset paper). Alternatively, a different Paper B evaluates a new prompting/evaluation protocol on the same dataset under a revised setup (result $\neqdataset paper$) (Park et al., 2021; Rangapur and Rangapur, 2024).
Dataset Source Paper Retrieval
@@
To compile faithful descriptions for each record, we first collect its sources. LLMEvalDB (Park et al., 2025) provides the result papers. When the result paper differs from the dataset paper, we automatically recover the latter by inferring the citation with an LLM, resolving it to an arXiv ID, downloading the PDF, and verifying the dataset name/title. We then run a rule-based check to confirm that the PDF explicitly mentions the dataset (string match).
Extracting Input Description
@@
Given the sources for the result and dataset papers, an extractor LLM produces a schema-aligned, identity-masked description $x_{i}$ that summarizes the task (problem definition and input/label format), data collection (provenance, domain, collection protocol), evaluation (metric and protocol), difficulty cues (number of classes, long context, reasoning depth), the relevant subset/split, the prompting strategy (zero/few-shot, CoT), and any other record-specific settings (context-window, tool/retrieval use), when reported; see the Appendix E for illustrative examples.
@@
When the retrieved sources do not actually cover the experimental record (e.g., due to retrieval errors from 4), we discard the case by having the extractor return failed, thereby correcting residual retrieval mistakes. For the remaining records, we remove dataset names, numerical results, and summary statistics to produce faithful, self-contained, anonymized.
Data Statistics and Quality Control
@@
We construct a bounded yet diverse corpus via stratified sampling over source papers, enforcing caps per paper and per dataset plus per-source quotas to prevent any single source from dominating. This yields 2,290 instances of records spanning distinct 1,019 dataset papers and 1,022 unique result papers, covering many datasets and configurations rather than a single benchmark (Ye et al., 2023). Because Precog is LLM-assisted, we further audited 30 random descriptions under a fixed protocol—checking anonymization (binary: no dataset names or performance numbers), schema coverage (5-point), and source support vs. the PDFs (5-point). The results (Table 2) show well-anonymized, accurate, and comprehensively grounded descriptions suitable for text-only performance forecasting.